OpenStudy (anonymous):

Determine the type of boundary line and shading for the graph of the inequality -5x - y less than or greater to -2

OpenStudy (anonymous):

Cheese

OpenStudy (anonymous):

Really -_-

OpenStudy (anonymous):

less than or greater to or less than or equal to

OpenStudy (anonymous):

<=

OpenStudy (anonymous):

okay plot the graph for -5x-y=-2

OpenStudy (anonymous):

\[-5x - y \le -2\]

OpenStudy (anonymous):

lol i am telling you to plot graph for -5x-y=-2 i wil tell the rest later

OpenStudy (anonymous):

ok hold on i will tell

OpenStudy (anonymous):

 so its <= -2 which region do you think should be shaded

OpenStudy (anonymous):

above

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

if its greater than y inequality then its above if its less than y inequality then its below

OpenStudy (anonymous):

ohhh!

OpenStudy (anonymous):

yes of u solve the equation lil. you will get \[y \le 2-5x\]

OpenStudy (anonymous):

so its y less than inequality

OpenStudy (anonymous):

 this will be it

OpenStudy (anonymous):

@electrokid m correct right?

OpenStudy (anonymous):

yes that looks like $$y\le2-5x$$

OpenStudy (anonymous):

right then :) there you go @ahoward79

OpenStudy (anonymous):

Dashed line with shading on the side that includes the origin. Solid line with shading on the side that does not include the origin. Dashed line with shading on the side that does not include the origin. Solid line with shading on the side that includes the origin.

OpenStudy (anonymous):

So B or D?

OpenStudy (anonymous):

what is this? is there a diagram in your problem?

OpenStudy (anonymous):

dashed line =? solid line = ? and it includes origin

OpenStudy (anonymous):

its D maybe you mean by solid line if line is included in graph yes line is included